# Effects of Large Optical Depth on CD Hi  21-cm Non-Gaussian Signal

### v1

@@ 3.6 Impact of large optical depth on bispectrum
The right columns of Figures 5 and 6 show the corresponding percentage change in the exact bispectra compared to the approximated bispectra calculated as, $\frac{\Delta^{3}_{\rm approx}-\Delta^{3}_{\rm exact}}{\Delta^{3}_{\rm exact}}\times 100\%$ as a function of redshift. We observe that, for most of the CD period, the magnitude of the exact bispectrum is suppressed compared to the approximate one. This suppression is mainly caused by the term $-\tau^{2}/2$ in the expression for the differential brightness temperature (equation 5), which reduces the value of $\delta T_{\rm b}$.
@@
Since $\tau^{2}\propto\rho^{2}_{\rm HI}/T^{2}_{S}$, the suppression is not uniform across the field. It is stronger in regions with a higher neutral hydrogen density and lower spin temperature. As a result, the $\tau^{2}/2$ term affects the non-Gaussian signal (i.e., bispectrum and skewness) more than the global or Gaussian component of the signal. For example, the percentage change in the bispectrum is found to be $\sim 10\%$ at redshift $z=18$ across all models and triangle configurations. As redshift decreases, this change increases, reaches a maximum at an intermediate redshift, and then declines again. This trend can be explained if we look at the redshift evolution of Hi 21-cm optical depth. As shown in Fig.1, the average optical depth increases, peaks at a certain redshift, and then decreases at lower redshifts. The percentage difference in the bispectrum becomes largest when the optical depth is highest, consistent with what is observed in the right columns of Figs. 5 and 6. We find that the percentage change in the bispectrum increases up to $\sim 250\%$ at redshifts $z\approx 14$ and $\approx 12$ for $f_{\rm x}=0.1$, for both the equilateral and squeezed triangles at $k_{1}=0.16\,{\rm Mpc}^{-1}$. A similar trend is observed on smaller scales, e.g., at $k_{1}=0.8\,{\rm Mpc}^{-1}$ (see Fig. 6) where the change reaches up to $\sim 300\%$. This large change is due to the fact that both the IGM kinetic temperature and hence the spin temperature remain very low in this case because heating is highly inefficient. In models with moderate heating ($f_{\rm x}=46.4$), the percentage change is smaller, typically lies between $\sim 10\%$ to $\sim 200\%$. This is because heating is more efficient than in the low $f_{\rm x}$ case. This results in higher IGM and spin temperatures. As a result, the Hi optical depth becomes lower which reduces the impact of higher-order terms in $\tau$. For models with very efficient heating ($f_{\rm x}=1000$), the IGM and spin temperatures are large from the beginning of the CD. This makes the contribution of higher-order terms less significant. Consequently, the change in the bispectrum is small, around $10\%$, in these cases. However, we find that the change reaches up to $\sim 200\%$ at a few redshifts. We also find that the percentage change can sometimes be negative. This typically occurs when the bispectrum transitions from negative to positive or vice versa.
@@ 4 Impact on power spectrum
Figure 7: This shows the redshift evolution of the spherically averaged dimensionless power spectrum for the approximated and exact $\delta T_{b}$, along with the percentage difference between them, at scales $\rm k=0.16\ Mpc^{-1}$ (upper two rows) and $\rm k=0.8\ Mpc^{-1}$ (lower two rows). In the first and third rows, the blue solid and dashed lines represent the power spectrum of the approximated and exact $\delta T_{b}$, respectively, for the $M_{\rm h,min}=10^{9}\ {\rm M}_{\odot}$ model, whereas the red solid and dashed lines represent the same for the $M_{\rm h,min}=10^{10}\ {\rm M}_{\odot}$ model. In the second and fourth rows, the blue and red solid lines show the corrsponding percentage differences between the approximated and exact cases for the $M_{\rm h,min}=10^{9}\ {\rm M}_{\odot}$ and $M_{\rm h,min}=10^{10}\ {\rm M}_{\odot}$ models, respectively.
@@
We can see, for $k=0.16$ Mpc-1, that at the beginning of the CD power spectrum increases as the $\rm{Ly{\alpha}}$ radiation starts coupling the spin temperature with the kinetic temperature. However, as the average $\tau$ is small at the beginning of CD, the change in the power spectrum between approximated and exact cases is small. As time progresses, $\rm{Ly{\alpha}}$ coupling becomes a dominating astrophysical event and as $\rm{Ly{\alpha}}$ radiation spreads throughout the IGM, the average $\tau$ becomes maximum around that redshift. Therefore, the change in the power spectrum is also maximum at that redshift. After that, when X-ray radiation starts heating the IGM, the average $\tau$ starts decreasing, which also decreases the change in the power spectrum. One notable difference here is that power spectra do not show any sign flip which can discriminate between these two different astrophysical events, whereas the bispectrum evolution shows a sign flip between these events (see Fig.5). Finally, when IGM gets almost entirely heated the power spectrum and its change both decreases.
@@ 5 Importance of different Higher-Order Terms of $\tau$ in Bispectrum Estimation
Figure 8: This shows the percentage differences between the exact and approximated bispectra of $\delta T_{\rm b}$ for the model with $f_{\rm x}=0.1$ and $M_{\rm h,min}=10^{9}\ {\rm M}_{\odot}$, at redshifts $z=12.3$, $14$, and $15.6$. The approximated $\delta T_{\rm b}$ includes contributions up to the first (blue), second (green), and third order (red) terms of $\tau$. The left and right panels correspond to the equilateral triangle configuration and the squeezed limit, respectively.
@@ 6 Summary and Discussions
The Hi 21-cm optical depth, which is proportional to the neutral hydrogen density and inversely proportional to the spin temperature, can become quite large during the Cosmic Dawn (CD). This is because the spin temperature is expected to remain very low before X-ray heating begins. Our simulations show that, in some CD models, the optical depth can reach values as high as $\approx 0.55$, with nearly $50\%$ of the simulation pixels showing optical depths greater than $0.1$. Therefore, the commonly used approximation that the optical depth is much less than one ( $\tau<<1$) which simplifies the expression $[1-\exp(-\tau)]\approx\tau$ may not hold during this epoch.
@@
We find that $\delta T_{\rm b}$ is significantly suppressed when the full expression with $[1-\exp(-\tau)]$ is used, particularly in regions with high neutral hydrogen density and low spin temperature. This suppression is not uniform across the volume but is stronger in regions where the ratio $\rho_{\rm HI}/T_{s}$ is higher. As a result, a new kind of non-Gaussianity (due to higher order terms of $\tau$ ) emerges with a sign opposite to that of the existing bispectrum. In the model with $M_{\rm h,min}=10^{9}\,{\rm M}_{\odot}$ and $f_{\rm x}=0.1$, the skewness changes by up to $\sim 500\%$. For the same halo mass, the change is about $40\%$ and $60\%$ for $f_{\rm x}=46.4$ and $f_{\rm x}=1000$, respectively. For a higher halo mass threshold of $M_{\rm h,min}=10^{10}\,{\rm M}_{\odot}$, the skewness changes by approximately $75\%$, $25\%$, and $20\%$ for $f_{\rm x}=0.1$, $46.4$, and $1000$, respectively.
@@ 7 Acknowledgments
IN acknowledges financial support from the Swami Vivekananda Merit-cum-Means Scholarship (SVMCM), Govt. of West Bengal, India. IN also acknowledges support from the DST-WISE Fellowship (DST/WISE-PhD/PM/2023/104), Govt. of India. KKD acknowledges financial support from ANRF (Govt. of India) through a SERB MATRICS scheme (MTR/2021/000384). The authors also acknowledge the use of computational resources provided by the Inter-University Centre for Astronomy and Astrophysics (IUCAA), Pune. RG acknowledges support from SERB, DST Ramanujan Fellowship no. RJF/2022/000141.

### v2

@@ 3.6 Impact of large optical depth on bispectrum
The right columns of Figures 5 and 6 show the corresponding percentage change in the exact bispectra compared to the approximated bispectra calculated as, $\frac{\Delta^{3}_{\rm approx}-\Delta^{3}_{\rm exact}}{\Delta^{3}_{\rm exact}}\times 100\%$ as a function of redshift. We observe that, for most of the CD period, the magnitude of the exact bispectrum is suppressed compared to the approximate one. This suppression is mainly caused by the term $\tau^{2}/2$ in the expression for the differential brightness temperature (equation 5), which reduces the value of $\delta T_{\rm b}$.
@@
Since $\tau^{2}\propto\rho^{2}_{\rm HI}/T^{2}_{S}$, the suppression is not uniform across the field. It is stronger in regions with a higher neutral hydrogen density and lower spin temperature. As a result, the $\tau^{2}/2$ term affects the non-Gaussian signal (i.e., the bispectrum and skewness) more strongly than it does the global 21-cm signal or the power spectrum. For example, the percentage change in the bispectrum is found to be $\sim 10\%$ at redshift $z=18$ across all models and triangle configurations. As redshift decreases, this change increases, reaches a maximum at an intermediate redshift, and then declines again. This trend can be explained if we look at the redshift evolution of Hi 21-cm optical depth. As shown in Fig.1, the average optical depth increases, peaks at a certain redshift, and then decreases at lower redshifts. The percentage difference in the bispectrum becomes largest when the optical depth is highest, consistent with what is observed in the right columns of Figs. 5 and 6. We find that the percentage change in the bispectrum increases up to $\sim 250\%$ at redshifts $z\approx 14$ and $\approx 12$ for $f_{\rm x}=0.1$, for both the equilateral and squeezed triangles at $k_{1}=0.16\,{\rm Mpc}^{-1}$. A similar trend is observed on smaller scales, e.g., at $k_{1}=0.8\,{\rm Mpc}^{-1}$ (see Fig. 6) where the change reaches up to $\sim 300\%$. This large change is due to the fact that both the IGM kinetic temperature and hence the spin temperature remain very low in this case because heating is highly inefficient. In models with moderate heating ($f_{\rm x}=46.4$), the percentage change is smaller, typically lies between $\sim 10\%$ to $\sim 200\%$. This is because heating is more efficient than in the low $f_{\rm x}$ case. This results in higher IGM and spin temperatures. As a result, the Hi optical depth becomes lower which reduces the impact of higher-order terms in $\tau$. For models with very efficient heating ($f_{\rm x}=1000$), the IGM and spin temperatures are large from the beginning of the CD. This makes the contribution of higher-order terms less significant. Consequently, the change in the bispectrum is small, around $10\%$, in these cases. However, we find that the change reaches up to $\sim 200\%$ at a few redshifts. We also find that the percentage change can sometimes be negative. This typically occurs when the bispectrum transitions from negative to positive or vice versa.
@@ 4 Impact on power spectrum
Figure 7: Redshift evolution of the spherically averaged dimensionless power spectrum for the approximated and exact $\delta T_{b}$, along with the percentage difference between them, at scales $\rm k=0.16\ Mpc^{-1}$ (upper two rows) and $\rm k=0.8\ Mpc^{-1}$ (lower two rows). In the first and third rows, the blue solid and dashed lines represent the power spectrum of the approximated and exact $\delta T_{b}$, respectively, for the $M_{\rm h,min}=10^{9}\ {\rm M}_{\odot}$ model, whereas the red solid and dashed lines represent the same for the $M_{\rm h,min}=10^{10}\ {\rm M}_{\odot}$ model. In the second and fourth rows, the blue and red solid lines show the corrsponding percentage differences between the approximated and exact cases for the $M_{\rm h,min}=10^{9}\ {\rm M}_{\odot}$ and $M_{\rm h,min}=10^{10}\ {\rm M}_{\odot}$ models, respectively.
@@
We can see, for $k=0.16$ Mpc-1, that at the beginning of the CD power spectrum increases as the $\rm{Ly{\alpha}}$ radiation starts coupling the spin temperature with the kinetic temperature. However, as the average $\tau$ is small at the beginning of CD, the change in the power spectrum between approximated and exact cases is small. As time progresses, $\rm{Ly{\alpha}}$ coupling becomes a dominating astrophysical event and as $\rm{Ly{\alpha}}$ radiation spreads throughout the IGM, the average $\tau$ becomes maximum around that redshift. Therefore, the change in the power spectrum is also maximum at that redshift. After that, when X-ray radiation starts heating the IGM, the average $\tau$ starts decreasing, which also decreases the change in the power spectrum. Finally, when IGM gets almost entirely heated the power spectrum and its change both decreases.
@@ 5 Importance of different Higher-Order Terms of $\tau$ in Bispectrum Estimation
Figure 8: Percentage differences between the exact and approximated bispectra of $\delta T_{\rm b}$ for the model with $f_{\rm x}=0.1$ and $M_{\rm h,min}=10^{9}\ {\rm M}_{\odot}$, at redshifts $z=12.3$, $14$, and $15.6$. The approximated $\delta T_{\rm b}$ includes contributions up to the first (blue), second (green), and third order (red) terms of $\tau$. The left and right panels correspond to the equilateral triangle configuration and the squeezed limit, respectively.
@@ 6 Summary and Discussions
The Hi 21-cm optical depth, which is proportional to the neutral hydrogen density and inversely proportional to the spin temperature, can become quite large during the Cosmic Dawn (CD). This is because the spin temperature is expected to remain very low before X-ray heating begins. Our simulations show that, in some CD models, the optical depth can reach values as high as $\approx 0.55$, with nearly $50\%$ of the simulation pixels showing optical depths greater than $0.1$. Therefore, the commonly used approximation that the optical depth is much less than one ( $\tau<<1$) which simplifies the expression $1-\exp(-\tau)]\approx\tau$ may not hold during this epoch.
@@
We find that $\delta T_{\rm b}$ is significantly suppressed when the full expression with $1-\exp(-\tau)]$ is used, particularly in regions with high neutral hydrogen density and low spin temperature. This suppression is not uniform across the volume but is stronger in regions where the ratio $\rho_{\rm HI}/T_{s}$ is higher. As a result, an additional non-Gaussian component arising from higher-order terms of $\tau$ emerges, with a sign opposite to that of the existing bispectrum. In the model with $M_{\rm h,min}=10^{9}\,{\rm M}_{\odot}$ and $f_{\rm x}=0.1$, the skewness changes by up to $\sim 500\%$. For the same halo mass, the change is about $40\%$ and $60\%$ for $f_{\rm x}=46.4$ and $f_{\rm x}=1000$, respectively. For a higher halo mass threshold of $M_{\rm h,min}=10^{10}\,{\rm M}_{\odot}$, the skewness changes by approximately $75\%$, $25\%$, and $20\%$ for $f_{\rm x}=0.1$, $46.4$, and $1000$, respectively.
@@
Our analysis suggests that for non-gaussian study of $\delta T_{\rm b}$ signal we need to incorporate higher order terms of $\tau$ along with the linear order term. It demonstrates that higher order terms of $\tau$ can have considerable contribution to the non-Gaussian statistics during early stages of CD. However, we can still use the approximated form of $\delta T_{\rm b}$ (eq.4) towards the later stages of CD and during EoR when the spin temperature is high i.e. $\tau$ is low.
@@ 7 Acknowledgments
IN acknowledges financial support from the Swami Vivekananda Merit-cum-Means Scholarship (SVMCM), Govt. of West Bengal, India. IN also acknowledges support from the DST-WISE Fellowship (DST/WISE-PhD/PM/2023/104), Govt. of India. KKD acknowledges financial support from ANRF (Govt. of India) through a SERB MATRICS scheme (MTR/2021/000384). The authors also acknowledge the use of computational resources provided by the Inter-University Centre for Astronomy and Astrophysics (IUCAA), Pune. AKS acknowledges support from National Science Foundation (grant no. 2206602). RG acknowledges support from SERB, DST Ramanujan Fellowship no. RJF/2022/000141. AM acknowledges financial support from Council of Scientific and Industrial Research (CSIR) via CSIR-SRF fellowships under grant no. 09/0096(13611)/2022-EMR-I. MK acknowledges financial support from the foundations Carl Tryggers stiftelse för vetenskaplig forskning (grant agreement no. CTS 21:1376).